estion 20 of 25 > Assign formal charges to cach atom in the three resonance forms of N,O. N-N=0: :N=N- N=n=; Answer Bank
Answer #1

IN – NS:0 i We know that formal Chorce = No. of Valence electrina armon-honded efecto y I fino of bonds) Mr - for pst Nitrogi- MilliN - N= 0 for Zizoi FC = 5-(4+2) = - Y PC a 5-c0+4)=1 1FC6-C472-0 | Scanned with CamScanner